# Nav-$E$: Navigation-Guided Early Exiting for Efficient Vision-Language Models in Autonomous Driving

### v1

@@ Abstract
Vision-Language Models (VLMs) are increasingly applied in autonomous driving for unified perception and reasoning, but high inference latency hinders real-time deployment. Early-exit reduces latency by terminating inference at intermediate layers, yet its task-dependent nature limits generalization across diverse scenarios. We observe that this limitation aligns with autonomous driving: navigation systems can anticipate upcoming contexts (e.g., intersections, traffic lights), indicating which tasks will be required. We propose Nav-$E$, a navigation-guided early-exit framework that precomputes task-specific exit layers offline and dynamically applies them online based on navigation priors. Experiments on CODA, Waymo, and BOSCH show that Nav-$E$ achieves accuracy comparable to full inference while reducing latency by up to 63.9%. Real-vehicle integration with Autoware Universe further demonstrates reduced inference latency (600 ms to 300 ms), supporting faster decision-making in complex scenarios. These results suggest that coupling navigation foresight with early-exit offers a viable path toward efficient deployment of large models in autonomous systems. Code and data are available at our anonymous repository: https://anonymous.4open.science/r/Nav-$E$-BBC4
@@ I INTRODUCTION
Figure 1: Full-layer inference processes all transformer layers, while generic $E$ methods stop once predictions stabilize. With navigation priors, VLMs in autonomous driving can adopt more aggressive exits, avoiding redundant computation while preserving accuracy.
Among sparsification techniques, early exit ($E$) is particularly attractive, as it halts inference once predictions stabilize [9, 10, 11]. As shown in Fig. 2, correct recognition often emerges in early layers, while later computation adds little value. $E$ therefore offers strong potential for reducing latency in driving systems, where avoiding unnecessary inference directly improves response time and efficiency.
However, $E$ strategies often lack generalization[34]. Prior work shows that in-domain or task-specific $E$ can deliver substantial computation savings, whereas out-of-domain or general-purpose benchmarks (e.g., GLUE) see limited benefits [9, 11, 12]. As illustrated in Fig. 1, confidence dynamics vary widely across domains, restricting aggressive exits in general-purpose settings. In contrast, for narrow tasks such as text classification or domain-focused applications, $E$ can skip up to half of the layers with negligible accuracy loss [9, 12]. This disparity suggests that $E$ is highly effective in specialized domains but unreliable when applied broadly.
Interestingly, this limitation aligns naturally with autonomous driving. Navigation systems can anticipate upcoming contexts—such as intersections, traffic lights, or pedestrian zones—thus providing strong priors about the perception or decision-making tasks that will soon be required. As illustrated in Fig. 3, these priors enable task-specific $E$ decisions, allowing VLMs to exit earlier without compromising accuracy.
In this work, we extend $E$ from unimodal or NLP models to large VLMs for autonomous driving. We exploit the determinism of navigation priors to adopt an aggressive statistical strategy for selecting exit layers, and design a dynamically switchable $E$ configuration that adapts to driving contexts in real time. We evaluate Nav-$E$ on large-scale datasets (Waymo [14], CODA [15], and Bosch) and validate it on a real vehicle integrated with Autoware.Universe. By monitoring HD-map ROS 2 topics to trigger task switches, Nav-$E$ achieves context-aware early exiting in practice, reducing latency while maintaining or improving accuracy.
@@
In summary, the challenge is that conventional $E$ strategies cannot generalize across the diverse and dynamic tasks in autonomous driving. Our contribution is to adapt $E$ to large VLMs and leverage deterministic navigation priors as a strength: they decompose driving into predictable sub-tasks, enabling reliable, task-aware acceleration in practice.
@@
  • We conduct extensive experiments across diverse datasets and models: the Waymo dataset for common perception tasks, the CODA dataset for animal detection and corner-case recognition, and the Bosch dataset for traffic-light decision making. Results consistently show that Nav-$E$ reduces inference latency and resource consumption while simultaneously improving accuracy.
  • We further validate Nav-$E$ in real-world deployment by integrating it into an autonomous vehicle running Autoware.Universe. The real-vehicle experiments confirm that navigation-guided early exit can operate reliably under practical driving conditions, achieving both efficiency gains and accuracy improvements.
@@ II-B Early Exit and Adaptive Inference
Early exit ($E$) reduces inference cost by halting computation once confidence is sufficient, and has been validated across BERT-like backbones and downstream tasks [9, 10, 11]. Post-deployment acceleration via block bypass and divergence criteria further improves practicality [12, 13]. For multimodal and generative settings, synchronized/robust $E$ and confidence-adaptive decoding have also been explored [18, 17]. Closest to our setting, retrieval-augmented early-exiting (RAEE) shows that instance-wise priors can guide exit-layer selection without re-training the backbone [16]. Recent work has also explored self-distillation for stabilizing exits [29], entropy-based criteria for adaptive decoding [30], and layer-wise calibration methods for robust $E$ under distribution shift [31]. However, a persistent limitation is that $E$ often depends on task/domain-specific configurations, hindering generalization. Our work embraces this constraint and turns it into an advantage by leveraging high-certainty navigation priors in autonomous driving to choose aggressive (earlier) exits safely.
@@ III METHODOLOGY
The complete Nav-$E$ pipeline consists of three stages (Fig. 4). First, in the offline profiling stage, we run the VLM across the training dataset and record predictions at every intermediate layer. Based on accuracy statistics, we identify the earliest valid exit layer for each scenario-specific task using an aggressive statistical selection strategy. Second, in the online inference stage, the system dynamically switches between these precomputed exit configurations according to navigation-derived scene contexts, ensuring that the most suitable exit layer is activated for each driving condition. Finally, in the real-vehicle deployment stage, Nav-$E$ is integrated into the Autoware.Universe stack, where HD map updates from ROS2 topics trigger real-time task-specific configuration switching.
@@ III-B Problem Formulation
Early-exit ($E$) allows inference to terminate at intermediate layers when a confidence criterion is satisfied [9, 10, 11]:
@@
Let $\text{Acc}(l)$ and $\text{Lat}(l)$ denote the accuracy and latency when exiting at layer $l$. Conventional $E$ methods adopt conservative thresholds to preserve accuracy, often resulting in late exits and limited efficiency. In contrast, we leverage the determinism of autonomous driving tasks to employ an aggressive statistical strategy, which selects the earliest valid exit layer by solving:
@@ III-C Offline Profiling: Aggressive Statistical Layer Selection
where NN is the number of training samples, $h_{l}$ is the hidden representation at layer $l$, $g(\cdot)$ is the shared classification head, and $y_{i}$ is the ground truth label.
@@ III-D Online Inference: Dynamic Switching for Autonomous Driving
After the optimal exit layers are profiled offline for different tasks, Nav-$E$ enables online adaptation through dynamic switching. In practice, navigation signals in autonomous driving provide reliable foresight of upcoming contexts (e.g., intersections, traffic lights, or pedestrian zones), as illustrated in Fig. 3. Leveraging this information, the system activates the corresponding pre-computed early-exit configuration for the predicted scene.
@@
To validate the practicality of our framework, we integrated Nav-$E$ into a real autonomous vehicle running the Autoware.Universe platform. The vehicle is equipped with six high-resolution cameras and GPU computing units onboard (RTX 3090), allowing direct deployment of VLM-based perception.
A key feature of Autoware.Universe is its access to high-definition (HD) maps and navigation signals via ROS2 topics. We subscribe to the HD map topic(/map/vector_map) to monitor upcoming driving contexts such as intersections, crosswalks, or traffic light regions. These navigation cues are then used to trigger dynamic switching of early-exit configurations: when the HD map indicates a new scene type $c_{k}$, the system immediately loads the corresponding offline-profiled exit layer $l^{\ast}_{c_{k}}$. This integration enables Nav-$E$ to operate seamlessly alongside existing perception and planning modules without requiring modification of the core Autoware stack. As shown in Fig. 5
@@ IV-A1 Datasets
We evaluate Nav-$E$ on three representative autonomous driving datasets. Waymo Open Dataset [14] provides large-scale real-world perception data with common driving scenarios, including vehicles, cyclists, and pedestrians. We adopt it for evaluating standard perception tasks such as vehicle and pedestrian recognition. CODA Dataset [15] focuses on challenging corner cases, covering rare categories such as animals, special vehicles, and unusual obstacles. CODA is particularly valuable for testing the robustness of VLMs under safety-critical conditions and highlights their advantages in recognizing uncommon objects. Bosch Small Traffic Lights Dataset is designed for traffic-light state recognition, containing diverse traffic light instances under varying illumination and occlusion conditions. We leverage this dataset to evaluate the decision-critical capability of Nav-$E$ in traffic signal understanding.
@@ IV-A2 VLMs Configuration
We adopt three representative Vision-Language Models (VLMs) to evaluate Nav-$E$ under different model capacities. LLaVA-v1.6-Vicuna-7B [6] is a large-scale VLM with stronger reasoning ability, serving as a testbed for evaluating the scalability of early-exit strategies on high-capacity models. Together, these models cover a spectrum from resource-efficient to large-scale VLMs, providing a comprehensive evaluation of Nav-$E$ across different computational regimes. DeepSeek-VL2-Tiny and DeepSeek-VL2-Small [5] are lightweight variants designed for efficient multimodal reasoning, making them suitable for real-time perception benchmarks.
@@ IV-A3 Task Selection
We focus on two types of tasks to evaluate Nav-$E$. Perception tasks. We select sidewalk pedestrian detection, road vehicle detection, road cyclist recognition, and animal recognition. On the Waymo Open Dataset, pedestrian and vehicle categories serve as canonical benchmarks for standard perception, reflecting how VLMs perform under typical urban driving conditions. On the CODA dataset, however, the same categories (pedestrian and vehicle) are significantly more challenging: as highlighted in the CODA paper [15], small perception models often fail to identify these objects under rare, occluded, or unusual conditions. This setting allows us to emphasize the comparative advantage of VLMs and the additional efficiency gains brought by early-exit. For animal recognition, CODA provides the most distinctive scenario: animals are rarely included in traditional autonomous driving datasets, making this an out-of-domain task for small perception models. VLMs, in contrast, exhibit stronger generalization to such categories, and our early-exit strategy highlights how efficiency can be preserved even in these rare but safety-critical cases. Decision task, We additionally evaluate traffic-light state recognition using the Bosch dataset. This task is safety-critical and directly governs driving decisions. Moreover, traffic-light understanding is a canonical case where VLMs can provide semantic guidance to complement smaller perception models, making it a natural testbed for evaluating decision-oriented capabilities. By combining both perception and decision tasks, our evaluation spans standard conditions, corner cases, and out-of-domain scenarios, demonstrating that navigation-guided early-exit not only improves efficiency but also amplifies the unique advantages of VLMs in autonomous driving.
@@ IV-C Perception Tasks
We evaluate perception tasks including pedestrians, cyclists, vehicles and animals. Table II summarizes both accuracy and latency on the CODA dataset. Compared with full-layer inference, our method (Nav-$E$) consistently improves accuracy across categories while also reducing inference time. This demonstrates that aggressive early-exiting not only avoids redundant computation but also mitigates over-inference, leading to more reliable intermediate-layer representations.
The results on CODA and WAYMO datasets highlight several important trends. First, for standard perception tasks such as pedestrian and vehicle recognition in WAYMO, Nav-$E$ significantly improves both accuracy and latency: DS-Tiny improves from 81.22% to 94.62% accuracy with a 22.0% latency reduction, and LLaVA-7B rises from 67.33% to 99.93% accuracy while reducing runtime by more than half. Second, for corner-case perception tasks in CODA, early-exit not only maintains but often surpasses full-layer performance. For example, vehicle and animal categories reach nearly perfect accuracy under Nav-$E$ while saving 15–60% runtime, showing that VLMs can stabilize predictions earlier without over-inference. Third, CODA results further demonstrate that VLMs complement smaller perception models: while DS-Tiny and DS-Small struggle with pedestrians (0% accuracy in full inference), their performance improves drastically under Nav-$E$ (e.g., pedestrian detection up to 52% in DS-SmallNav-$E$).
@@
LLaVA-7B (General-$E$) 67.11 87.38 91.18 37.68
LLaVA-7B (Nav-$E$) 100.00 99.98 100.00 76.93
@@
LLaVA-7B (General-$E$) 0.021 0.018 0.018 0.028
LLaVA-7B (Nav-$E$) 0.015 0.013 0.013 0.025
While early-exit is typically motivated by efficiency, our results reveal a critical phenomenon: over-thinking in full-layer inference can harm accuracy. We observe failure cases where mid-depth layers already yield correct predictions, but deeper layers drift toward errors due to noisy refinement or excessive reasoning. This aligns with prior findings in LLMs that deeper inference may induce hallucinations or unstable decision boundaries despite higher cost [26, 27]. For perception tasks such as vehicles or pedestrians, once high-confidence recognition emerges at intermediate layers, continued inference risks semantic confusion rather than improvement. By terminating at stable early layers, Nav-$E$ not only improves efficiency but also safeguards against over-inference, delivering faster and more reliable predictions in safety-critical driving scenarios.

### v2

# Nav-EE: Navigation-Guided Early Exiting for Efficient Vision-Language Models in Autonomous Driving

@@ Abstract
Vision-Language Models (VLMs) are increasingly applied in autonomous driving for unified perception and reasoning, but high inference latency hinders real-time deployment. Early-exit reduces latency by terminating inference at intermediate layers, yet its task-dependent nature limits generalization across diverse scenarios. We observe that this limitation aligns with autonomous driving: navigation systems can anticipate upcoming contexts (e.g., intersections, traffic lights), indicating which tasks will be required. We propose Nav-EE, a navigation-guided early-exit framework that precomputes task-specific exit layers offline and dynamically applies them online based on navigation priors. Experiments on CODA, Waymo, and BOSCH show that Nav-EE achieves accuracy comparable to full inference while reducing latency by up to 63.9%. Real-vehicle integration with Autoware Universe further demonstrates reduced inference latency (600 ms to 300 ms), supporting faster decision-making in complex scenarios. These results suggest that coupling navigation foresight with early-exit offers a viable path toward efficient deployment of large models in autonomous systems. Code and data are available at our anonymous repository: https://anonymous.4open.science/r/Nav-EE-BBC4
@@ I INTRODUCTION
Figure 1: Full-layer inference processes all transformer layers, while generic EE methods stop once predictions stabilize. With navigation priors, VLMs in autonomous driving can adopt more aggressive exits, avoiding redundant computation while preserving accuracy.
Among sparsification techniques, early exit (EE) is particularly attractive, as it halts inference once predictions stabilize [9, 10, 11]. As shown in Fig. 2, correct recognition often emerges in early layers, while later computation adds little value. EE therefore offers strong potential for reducing latency in driving systems, where avoiding unnecessary inference directly improves response time and efficiency.
However, EE strategies often lack generalization[34]. Prior work shows that in-domain or task-specific EE can deliver substantial computation savings, whereas out-of-domain or general-purpose benchmarks (e.g., GLUE) see limited benefits [9, 11, 12]. As illustrated in Fig. 1, confidence dynamics vary widely across domains, restricting aggressive exits in general-purpose settings. In contrast, for narrow tasks such as text classification or domain-focused applications, EE can skip up to half of the layers with negligible accuracy loss [9, 12]. This disparity suggests that EE is highly effective in specialized domains but unreliable when applied broadly.
Interestingly, this limitation aligns naturally with autonomous driving. Navigation systems can anticipate upcoming contexts—such as intersections, traffic lights, or pedestrian zones—thus providing strong priors about the perception or decision-making tasks that will soon be required. As illustrated in Fig. 3, these priors enable task-specific EE decisions, allowing VLMs to exit earlier without compromising accuracy.
In this work, we extend EE from unimodal or NLP models to large VLMs for autonomous driving. We exploit the determinism of navigation priors to adopt an aggressive statistical strategy for selecting exit layers, and design a dynamically switchable EE configuration that adapts to driving contexts in real time. We evaluate Nav-EE on large-scale datasets (Waymo [14], CODA [15], and Bosch) and validate it on a real vehicle integrated with Autoware.Universe. By monitoring HD-map ROS 2 topics to trigger task switches, Nav-EE achieves context-aware early exiting in practice, reducing latency while maintaining or improving accuracy.
@@
In summary, the challenge is that conventional EE strategies cannot generalize across the diverse and dynamic tasks in autonomous driving. Our contribution is to adapt EE to large VLMs and leverage deterministic navigation priors as a strength: they decompose driving into predictable sub-tasks, enabling reliable, task-aware acceleration in practice.
@@
  • We conduct extensive experiments across diverse datasets and models: the Waymo dataset for common perception tasks, the CODA dataset for animal detection and corner-case recognition, and the Bosch dataset for traffic-light decision making. Results consistently show that Nav-EE reduces inference latency and resource consumption while simultaneously improving accuracy.
  • We further validate Nav-EE in real-world deployment by integrating it into an autonomous vehicle running Autoware.Universe. The real-vehicle experiments confirm that navigation-guided early exit can operate reliably under practical driving conditions, achieving both efficiency gains and accuracy improvements.
@@ II-B Early Exit and Adaptive Inference
Early exit (EE) reduces inference cost by halting computation once confidence is sufficient, and has been validated across BERT-like backbones and downstream tasks [9, 10, 11]. Post-deployment acceleration via block bypass and divergence criteria further improves practicality [12, 13]. For multimodal and generative settings, synchronized/robust EE and confidence-adaptive decoding have also been explored [18, 17]. Closest to our setting, retrieval-augmented early-exiting (RAEE) shows that instance-wise priors can guide exit-layer selection without re-training the backbone [16]. Recent work has also explored self-distillation for stabilizing exits [29], entropy-based criteria for adaptive decoding [30], and layer-wise calibration methods for robust EE under distribution shift [31]. However, a persistent limitation is that EE often depends on task/domain-specific configurations, hindering generalization. Our work embraces this constraint and turns it into an advantage by leveraging high-certainty navigation priors in autonomous driving to choose aggressive (earlier) exits safely.
@@ III METHODOLOGY
The complete Nav-EE pipeline consists of three stages (Fig. 4). First, in the offline profiling stage, we run the VLM across the training dataset and record predictions at every intermediate layer. Based on accuracy statistics, we identify the earliest valid exit layer for each scenario-specific task using an aggressive statistical selection strategy. Second, in the online inference stage, the system dynamically switches between these precomputed exit configurations according to navigation-derived scene contexts, ensuring that the most suitable exit layer is activated for each driving condition. Finally, in the real-vehicle deployment stage, Nav-EE is integrated into the Autoware.Universe stack, where HD map updates from ROS2 topics trigger real-time task-specific configuration switching.
@@ III-B Problem Formulation
Early-exit (EE) allows inference to terminate at intermediate layers when a confidence criterion is satisfied [9, 10, 11]:
@@
Let $\text{Acc}(l)$ and $\text{Lat}(l)$ denote the accuracy and latency when exiting at layer $l$. Conventional EE methods adopt conservative thresholds to preserve accuracy, often resulting in late exits and limited efficiency. In contrast, we leverage the determinism of autonomous driving tasks to employ an aggressive statistical strategy, which selects the earliest valid exit layer by solving:
@@ III-C Offline Profiling: Aggressive Statistical Layer Selection
where $N$ is the number of training samples, $h_{l}$ is the hidden representation at layer $l$, $g(\cdot)$ is the shared classification head, and $y_{i}$ is the ground truth label.
@@ III-D Online Inference: Dynamic Switching for Autonomous Driving
After the optimal exit layers are profiled offline for different tasks, Nav-EE enables online adaptation through dynamic switching. In practice, navigation signals in autonomous driving provide reliable foresight of upcoming contexts (e.g., intersections, traffic lights, or pedestrian zones), as illustrated in Fig. 3. Leveraging this information, the system activates the corresponding pre-computed early-exit configuration for the predicted scene.
@@
To validate the practicality of our framework, we integrated Nav-EE into a real autonomous vehicle running the Autoware.Universe platform. The vehicle is equipped with six high-resolution cameras and GPU computing units onboard (RTX 3090), allowing direct deployment of VLM-based perception.
A key feature of Autoware.Universe is its access to high-definition (HD) maps and navigation signals via ROS2 topics. We subscribe to the HD map topic(/map/vector_map) to monitor upcoming driving contexts such as intersections, crosswalks, or traffic light regions. These navigation cues are then used to trigger dynamic switching of early-exit configurations: when the HD map indicates a new scene type $c_{k}$, the system immediately loads the corresponding offline-profiled exit layer $l^{\ast}_{c_{k}}$. This integration enables Nav-EE to operate seamlessly alongside existing perception and planning modules without requiring modification of the core Autoware stack. As shown in Fig. 5
@@ IV-A1 Datasets
We evaluate Nav-EE on three representative autonomous driving datasets. Waymo Open Dataset [14] provides large-scale real-world perception data with common driving scenarios, including vehicles, cyclists, and pedestrians. We adopt it for evaluating standard perception tasks such as vehicle and pedestrian recognition. CODA Dataset [15] focuses on challenging corner cases, covering rare categories such as animals, special vehicles, and unusual obstacles. CODA is particularly valuable for testing the robustness of VLMs under safety-critical conditions and highlights their advantages in recognizing uncommon objects. Bosch Small Traffic Lights Dataset is designed for traffic-light state recognition, containing diverse traffic light instances under varying illumination and occlusion conditions. We leverage this dataset to evaluate the decision-critical capability of Nav-EE in traffic signal understanding.
@@ IV-A2 VLMs Configuration
We adopt three representative Vision-Language Models (VLMs) to evaluate Nav-EE under different model capacities. LLaVA-v1.6-Vicuna-7B [6] is a large-scale VLM with stronger reasoning ability, serving as a testbed for evaluating the scalability of early-exit strategies on high-capacity models. Together, these models cover a spectrum from resource-efficient to large-scale VLMs, providing a comprehensive evaluation of Nav-EE across different computational regimes. DeepSeek-VL2-Tiny and DeepSeek-VL2-Small [5] are lightweight variants designed for efficient multimodal reasoning, making them suitable for real-time perception benchmarks.
@@ IV-A3 Task Selection
We focus on two types of tasks to evaluate Nav-EE. Perception tasks. We select sidewalk pedestrian detection, road vehicle detection, road cyclist recognition, and animal recognition. On the Waymo Open Dataset, pedestrian and vehicle categories serve as canonical benchmarks for standard perception, reflecting how VLMs perform under typical urban driving conditions. On the CODA dataset, however, the same categories (pedestrian and vehicle) are significantly more challenging: as highlighted in the CODA paper [15], small perception models often fail to identify these objects under rare, occluded, or unusual conditions. This setting allows us to emphasize the comparative advantage of VLMs and the additional efficiency gains brought by early-exit. For animal recognition, CODA provides the most distinctive scenario: animals are rarely included in traditional autonomous driving datasets, making this an out-of-domain task for small perception models. VLMs, in contrast, exhibit stronger generalization to such categories, and our early-exit strategy highlights how efficiency can be preserved even in these rare but safety-critical cases. Decision task, We additionally evaluate traffic-light state recognition using the Bosch dataset. This task is safety-critical and directly governs driving decisions. Moreover, traffic-light understanding is a canonical case where VLMs can provide semantic guidance to complement smaller perception models, making it a natural testbed for evaluating decision-oriented capabilities. By combining both perception and decision tasks, our evaluation spans standard conditions, corner cases, and out-of-domain scenarios, demonstrating that navigation-guided early-exit not only improves efficiency but also amplifies the unique advantages of VLMs in autonomous driving.
@@ IV-C Perception Tasks
We evaluate perception tasks including pedestrians, cyclists, vehicles and animals. Table II summarizes both accuracy and latency on the CODA dataset. Compared with full-layer inference, our method (Nav-EE) consistently improves accuracy across categories while also reducing inference time. This demonstrates that aggressive early-exiting not only avoids redundant computation but also mitigates over-inference, leading to more reliable intermediate-layer representations.
The results on CODA and WAYMO datasets highlight several important trends. First, for standard perception tasks such as pedestrian and vehicle recognition in WAYMO, Nav-EE significantly improves both accuracy and latency: DS-Tiny improves from 81.22% to 94.62% accuracy with a 22.0% latency reduction, and LLaVA-7B rises from 67.33% to 99.93% accuracy while reducing runtime by more than half. Second, for corner-case perception tasks in CODA, early-exit not only maintains but often surpasses full-layer performance. For example, vehicle and animal categories reach nearly perfect accuracy under Nav-EE while saving 15–60% runtime, showing that VLMs can stabilize predictions earlier without over-inference. Third, CODA results further demonstrate that VLMs complement smaller perception models: while DS-Tiny and DS-Small struggle with pedestrians (0% accuracy in full inference), their performance improves drastically under Nav-EE (e.g., pedestrian detection up to 52% in DS-SmallNav-EE).
@@
LLaVA-7B (General-EE) 67.11 87.38 91.18 37.68
LLaVA-7B (Nav-EE) 100.00 99.98 100.00 76.93
@@
LLaVA-7B (General-EE) 0.021 0.018 0.018 0.028
LLaVA-7B (Nav-EE) 0.015 0.013 0.013 0.025
While early-exit is typically motivated by efficiency, our results reveal a critical phenomenon: over-thinking in full-layer inference can harm accuracy. We observe failure cases where mid-depth layers already yield correct predictions, but deeper layers drift toward errors due to noisy refinement or excessive reasoning. This aligns with prior findings in LLMs that deeper inference may induce hallucinations or unstable decision boundaries despite higher cost [26, 27]. For perception tasks such as vehicles or pedestrians, once high-confidence recognition emerges at intermediate layers, continued inference risks semantic confusion rather than improvement. By terminating at stable early layers, Nav-EE not only improves efficiency but also safeguards against over-inference, delivering faster and more reliable predictions in safety-critical driving scenarios.